A ball of mass 0.1kg is thrown vertically upwards with initial velocity of 80m/s.Calculate the potential energy ( I) halfway up(11) at its maximum height. What is it kinetic energy as it leaves the ground?.
1
Expert's answer
2020-05-05T18:41:36-0400

Let's calculate the maximum height

"H=v^2\/2g=320m"

PE=mgh=0.1*10*320=320 J

at halfway up PE will be half of max = 320/2=160 J

KE will be equal to PEmax=320 J